OpenStudy (anonymous):

Write a formula for the slope of a line that is perpendicular to the line through the points (x1,y1) and (x2,y2)

OpenStudy (danjs):

The perpendicular slope is the negative reciprical of the given slope

OpenStudy (danjs):

m = \[\frac{ (y2-y1) }{ (x2-x1) }\]

OpenStudy (danjs):

perpendicular slope is: \[\frac{ -(x2-x1) }{ (y2-y1) }\]

OpenStudy (anonymous):

Sweet. Thanks just wanted to double check my answer

OpenStudy (danjs):

\[(y - y1) = \frac{ -(x2-x1) }{ y2-y1 } * (x - x1)\]

OpenStudy (anonymous):

Thanks!